OpenStudy (anonymous):

If f(x)= 2^x undergoes a transformation and becomes g(x)= -6f(x+1)-5 how would I graph that in a calculator? I forgot what to do with the f :/.

OpenStudy (anonymous):

If f(x) is 2^x then in the second we have f(x+1), so you plug (x-1) in for x in f(x), I think. f(x-1) = 2^(x-1)

OpenStudy (anonymous):

Sorry, it's +1

OpenStudy (anonymous):

^^

OpenStudy (anonymous):

Wait so what would -6f(x+1)+5 be written as on a graphing calculator?

OpenStudy (anonymous):

Well we're inputting the function f(x+1) so we'd have g(x)=-6f(x+1)+5 =-6(2^(x+1))+5

OpenStudy (anonymous):

If there aren't any brackets or anything missing, that's what we'd input.

OpenStudy (anonymous):

ohhh I get it, thanks! :)